OpenStudy (kobeni-chan):

Determine how many, what type, and find the roots for f(x) = x^3 - 5x^2 – 25x + 125.

OpenStudy (kobeni-chan):

using the fundamental theorem of algebra - will medal

OpenStudy (aum):

Factor x^3 - 5x^2 – 25x + 125 by grouping as follows: (x^3 - 5x^2) – (25x - 125)

OpenStudy (kobeni-chan):

ok, and then I find the gcf's?

OpenStudy (aum):

Yes. GCF of x^3 and -5x^2. Then GCF of -25x and +125.

OpenStudy (kobeni-chan):

ok so that'll be x^2(x-5) and 25(x+5) ?

OpenStudy (aum):

x^2(x-5) and -25(x-5)

OpenStudy (kobeni-chan):

oh ok. since the second part is negative, do I add the two?

OpenStudy (aum):

f(x) = x^3 - 5x^2 – 25x + 125 = x^2(x - 5) - 25(x - 5) = (x^2 - 25)(x - 5) = (x + 5)(x - 5)(x - 5) = (x + 5)(x - 5)^2

OpenStudy (aum):

To find the root, set f(x) = 0 and solve for x. (x+5)(x-5)^2 = 0 implies x = -5 or x = 5. The roots are: -5 and 5 (with multiplicity of 2).

OpenStudy (kobeni-chan):

ok, thank you! :)

OpenStudy (aum):

You are welcome. How many roots? Two What type of roots? Two real roots Find the roots: -5 and 5. 5 has a multiplicity of two.